# VitaGlyph: Vitalizing Artistic Typography with Flexible Dual-branch Diffusion Models

### v2

@@ 3.3.2 Controllable Compositional Generation
The overall Controllable Compositional Generation process is illustrated in Fig. 4. Since the sampling processes for the subject and surrounding are conducted independently, and the U-Net outputs are independent as well, directly merging them can result in a visually incoherent image, as seen in Fig.8 (b). In our method, we resolve this fragmentation issue by allowing the subject and surrounding diffusion processes to share the same unconditional noise $\epsilon^{\mathrm{uc}}_{t}$ during Controllable Compositional Generation. $\epsilon^{\mathrm{overall}}_{t}$ can be further harmonized as follows:
where $s$ represents the guidance scale.
@@ 4 Experiments
Figure 6: Visualization of VitaGlyph’s applications. The experiments show that our method can be applied to other languages and fonts. Additionally, VitaGlyph can be extended to generate with multiple customized concepts.
Methods Language CLIP Score $\uparrow$ GPT-4o(TIM) $\uparrow$ GPT-4o(OCR) $\uparrow$ FID Score $\downarrow$
WordArt Designer [11] Chinese 23.93 94.98 88.67 249.12
VitaGlyph 25.40 96.38 92.89 239.27
Word-as-image [17] English 29.01 71.17 80.52 255.28
VitaGlyph 29.16 75.98 81.96 247.17
@@ 4.1 Implementation Details
4.1.1 Benchmark
We collect two hundred Chinese words (including single and multiple characters) as an evaluation benchmark and ask ChatGPT to annotate subject and surrounding descriptions. Considering that Word-as-image [17] only generates English words, we also translate the collected benchmark into English. More benchmark details are provided in the supplementary materials.
4.1.2 Evaluation Metrics
We use CLIP [38] image-text similarity and FID [13] scores to assess prompt alignment and readability. To better align with human evaluators, we also employ GPT-4o [33] to determine whether the rendered images match the text descriptions (GPT-4o Text-Image Matching) and can be correctly recognized as input characters (GPT-4o OCR).
@@ 4.2 Quantitative Comparisons
As shown in Tab. 1, we conduct quantitative experiments on the benchmark mentioned above. For comparisons, we leverage WordArt Designer [11] for Chinese words and Word-as-image [17] for English version. Since Word-as-image can only generate black text with white backgrounds, we apply ControlNets on it for texture rendering, as done in prior work [11]. Since MetaDesigner has not open-sourced the code and we can only generate results one by one through the website demo, we have only conducted a qualitative comparison with it.
Our VitaGlyph outperforms WordArt Designer in all metrics, especially in OCR with a 4.22% improvement. We believe this is because, with the controllable and proper transformation of glyph geometry, our method maintains promising readability. Meanwhile, for English words, VitaGlyph also demonstrates strong generalization capabilities, especially in TIM with a 4.81% improvement. This further proves that our method can better express essential concept with elaborately designed subject transformations.
Since all methods’ output images are in color while the ground-truth images for FID evaluation are black and white, all methods result in high FID scores. However, we still outperform other baselines.
@@ 4.3 Qualitative Comparisons
We compare VitaGlyph with WordArt Designer and MetaDesigner, both of which are strong baselines for Chinese artistic typography. We also design English words for a qualitative comparison with Word-as-image. The input prompts of WordArt Designer, MetaDesigner and Word-as-image consist of the concatenation of the subject descriptions and the surrounding descriptions. The input prompts for the English word is manually designed to include Nashville’s specialties, such as “country music, guitar, fruit” and so on.
@@
In Fig. 5, we see that VitaGlyph renders input characters with higher quality and legibility than WordArt Designer and achieves results comparable to, or even better than, MetaDesigner. For example, in column $1$, VitaGlyph visualizes the concept of “tomato” more concisely, incorporating related elements like “fries.” However, the image from WordArt Designer suffers from structural damage, and that of MetaDesigner blends the subject and background, visibly harming readability. As for Word-as-Image, it excels advantage in structural transformation but lacks artistic flair as it can only generate black-and-white typography. In contrast, VitaGlyph properly generates flexible and imaginable artistic typography with the controllable transformation. More examples generated by our VitaGlyph can be found in Fig. 1 and supplementary materials.
4.4 User Study
Methods Text Align. Char. Align. Artistic
@@ 4.4 User Study
Table 2: User Preference Study. Thirty users participated in the survey, selecting images that best represent the concept (Text Align.), match the character (Char. Align.), and show the highest artistic mastery (Artistic).
@@
We conduct a user study to evaluate our method against WordArt Designer and Word-as-Image. For customized multi-concept generation, users are shown the generated artistic text from each method along with ten reference images of the character. They are then asked to select the image that most represents the concept, the image that best matches the character itself, and the image that is most aesthetically pleasing. The questionnaire contains 80 sets for each question, and 30 responses were collected in total. The statistical results in Tab. 2 show that users prefer the rendered images from VitaGlyph across all metrics.
4.5 Ablation Studies
[…]
TIM $\uparrow$ OCR $\uparrow$
w/o $\texttt{GDINO}_{\mathrm{rank}}$ 23.90 94.95 89.15 255.34
w/o SemTypo 23.53 95.27 92.98 241.58
w/o shared-$\epsilon_{T}$ 23.82 95.22 90.38 244.77
Ours 25.40 96.38 92.89 239.27
Table 3: Ablation of components in our method. We provide our ablation studies’ quantitative results with different modules removed.
4.5.1 Effect of Region Interpretation
To study the effect of our Region Interpretation from Sec.3.2.2, we conduct ablation experiments without $\texttt{GDINO}_{\mathrm{rank}}$. Specifically, we regard $I$ as $I_{\mathrm{sub}}$, while keeping the settings for SemTypo and CCG unchanged. We solely utilize a $\mathrm{controlnet}_{\mathrm{sub}}$ with $I^{*}_{\mathrm{sub}}$ as the additional condition. The input prompt is a concatenation of $P_{\mathrm{sub}}$ and $P_{\mathrm{surr}}$. The results are shown in Fig. 7 (a). Additionally, we also randomly select the subject region rather than utilize $\texttt{GDINO}_{\mathrm{rank}}$, with all other settings unchanged. The results are presented in Fig. 7 (b).
Fig. 7 shows that without Region Interpretation, VitaGlyph faces challenges: In Fig. 7 (a), the overall deformation of characters is abnormal and some key concepts, such as “fries” in “tomato” (3rd row), are missing. In Fig. 7 (b), unreasonable region selection makes it difficult to ensure both prompt alignment and readability simultaneously. For example, the structure of “lavender” is deformed (1st row) and the features of “cat” are not accurately generated (2nd row).
For quantitative results, comparing to rows 2 and 5 of Tab. 3, the CLIP score improves by 1.5%, and accuracy increases by 1.43% in GPT-4o(TIM) and 3.09% in GPT-4o(OCR). This is likely because Grounding-DINO identifies the part most resembling the subject prompt, making glyph geometry transformation easier and more natural.
4.5.2 Effect of Semantic Typography
Figure 8: Visualization of ablation studies on SemTypo and shared-$\boldsymbol{\epsilon_{T}}$. Input characters are shown at the top of the sub-figures. The different columns in sub-figures (a) and (b) display experimental results with/without SemTypo or shared-$\epsilon_{T}$.
Although Grounding-DINO identifies the part similar to the subject prompt $P_{\mathrm{sub}}$, the Semantic Typography is essential for structural transformation. Ablation experiments without this module show that images fail to capture the character’s essence or cause the Subject to appear fragmented in Fig. 8 (a). This is due to the gap between $I_{\mathrm{sub}}$ and real-world objects, which the Depth-to-image module helps bridge.
Based on the rows 3 and 5 from Tab. 3, we can conclude that VitaGlyph with Semantic Typography would enrich the semantic expression but sacrifice little readability.
4.5.3 Effect of shared-$\boldsymbol{\epsilon_{T}}$
As VitaGlyph employs two ControlNets to individually render subject and surrounding regions, we investigate the impact of shared initial noise (i.e., shared-$\epsilon_{T}$) on the final results. Fig. 8 (b) indicates that not using shared-$\epsilon_{T}$ leads to inconsistent appearance and unrealistic lighting. For example, in the 1st column of “rose” and “monkey”, the colors of subject and surrounding regions are visibly incoherent. Additionally, rows 4 and 5 of Tab. 3 show that using shared-$\epsilon_{T}$ improves prompt alignment and readability, consistent with visualization results.
4.5.4 The Strength of Added Noise in SemTypo
Figure 9: Ablation Study on D2I’s strength $\mathrm{DS}$ and ControlNet’s scale $\mathrm{CS}$. (a) shows that as $\mathrm{DS}$ increases, the deformation of $I^{*}_{\mathrm{sub}}$ grows during SemTypo. (b) demonstrates the generated image morphology aligns more closely with $I^{*}_{\mathrm{sub}}$ as $\mathrm{CS}$ increases.
Semantic Typography uses the SDEdit algorithm [28] based on depth-to-image diffusion model to adjust the geometry of subject region. As shown in Fig. 9 (a), as the strength of added noise ($\mathrm{DS}$) increases, more structural details are added to the subject region while remaining the black-and-white color scheme unchanged. When $\mathrm{DS}$ reaches $0.85$, the subject region contains enough details of the given concept, e.g., the ears and tails of cats.
4.5.5 The Conditioning Scale in ControlNet
We fix the $I^{*}_{\mathrm{sub}}$ and only adjust the $\mathrm{ControlNet}_{\mathrm{sub}}$’s scale $\mathrm{CS}$. From Fig. 9 (b), we observe that as the $\mathrm{CS}$ increases, the generated part of Subject becomes more clear and more similar to the $I^{*}_{\mathrm{sub}}$, shown in row 1. However, being too similar to the subject can lead to detail loss, as seen in the row 2 where the flower’s stamen disappears from $\mathrm{CS}=1.0$.
4.6 Applications on Fonts and Languages.
In first row of Fig. 6, we demonstrate VitaGlyph’s effectiveness across various fonts and languages, such as “Bold” in English, “Song Traditional” in Chinese and “Noto Sans” in Korean. VitaGlyph not only generates artistic glyph images but also preserves the original font styles.
Moreover, VitaGlyph extends from single to multiple customized subject concepts. As illustrated in Fig. 6, it designs subjects like “the Earth”, “compass” and “colorful doughnut” within the artistic typography of “Google”. This plays a crucial role in areas such as commercial advertising, brand identity, and corporate image design. More examples can be seen in Fig. 1 and supplementary materials.
5 Conclusion
This paper presents VitaGlyph, a novel approach to artistic typography that balances creativity and readability through controllable glyph transformation. By sdistinguishing between the subject and surrounding elements, VitaGlyph enhances visual representation while preserving the character’s core meaning. Our method makes significant contributions to artistic text generation. Future work will focus on further advancements in generative AI to enhance the adaptability of typography across diverse contexts and styles. VitaGlyph will significantly enrich artistic expression and enabling more personalized, dynamic designs.
F Benchmark Details
We collect two hundred Chinese words (including single and multiple characters) as evaluation benchmark, as shown in Fig .K. Every word we generate one hundred samples and we totally create twenty thousand images as a dataset. Considering that Word-as-image [17] only generates English words, we also translate the collected benchmark into an English version with fifty samples per word and ten thousand images as a dataset at all.
For words’ descriptions, we ask ChatGPT [32] to annotate subject and surrounding prompts. Since other method, Word-as-image [17], WordArt Designer [11] and MetaDesigner [12] can only leverage one input prompt, we concatenate the subject and surrounding prompts.
G Deployment Details
In phase Knowledge Acquisition, we use GPT3.5 as our Large Language Model and the output of the LLM is JSON format. In phase Regional Decomposition, we use Grounding-DINO with its hyper-parameters all default to the  [23]. After that, we select the bounding boxes with its confidence over 0.5 and its area proportion of the whole image between 0.4 and 0.6 and then choose the highest confidence as the subject part. In phase Typography Stylization, we use Stable Diffusion-1.5 as our base model. All generated images are $512\times 512$ for both Depth-to-image and CCG module. Though we haven’t used larger modules (e.g., SD-XL [35] and DALL-E 3 [5]) to generate higher quality images, our results still demonstrate a strong generative capability. It is easy to deployment our work with one Nvidia RTX 3060 12GB GPU, taking about 40 seconds to generate 1 image.
H Additional Results

### v3

@@ 3.3 Attentional Compositional Generation
Cross-branch Attention
Building on coarse-grained fusion rationale (Eq. 4), we replace simple noise-based blending (seen in supplementary materials) with attention-driven fusion. To empower the Surrounding branch to carry Subject semantics, we insert dedicated cross-branch attention layers into the the $\mathrm{ControlNet}_{\mathrm{surr}}$ rather than relying on weight mixing. Concretely, during the Surrounding denoising process, we extract query-key pairs from $\epsilon_{t}$ with projection matrices $W_{q}$ and $W_{k}$ as
$q^{\mathrm{surr}}_{t}=\mathcal{F}^{\mathrm{surr}}_{t}\cdot W_{q},k^{\mathrm{surr}}_{t}=\mathcal{F}^{\mathrm{surr}}_{t}\cdot W_{k},$ (5)
where $(q^{\mathrm{surr}}_{t},k^{\mathrm{surr}}_{t})$ represents the query-key pair calculated from intermediate feature map $\mathcal{F}^{\mathrm{surr}}_{t}$ at every layer of U-Net. Self-attention scores are generally calculated as $\mathcal{A}^{\mathrm{surr}}_{t}=q^{\mathrm{surr}}_{t}\cdot(k^{\mathrm{surr}}_{t})^{T}/\sqrt{d_{dim}}$. Instead, we compose these self-attention scores with cross-attention against Subject query-key pairs $(q^{\mathrm{sub}}_{t},k^{\mathrm{sub}}_{t}))$. As discussed in (Feng et al. 2024), self-attention scores naturally encodes significant semantic cues. To leverage this for cross-region interaction, we synthesize the cross-branch key with a hyperparameter $\alpha$ [^2] We set the $\alpha$ as 0.5. as
$k^{\mathrm{cross}}_{t}=\alpha\cdot k^{\mathrm{surr}}_{t}+(1-\alpha)\cdot k^{\mathrm{sub}}_{t}$ (6)
and compute the cross-branch attention:
$\mathcal{A}^{\mathrm{cross}}_{t}=q^{\mathrm{surr}}_{t}\cdot(k^{\mathrm{cross}}_{t})^{T}/\sqrt{d_{dim}},$ (7)
seamlessly replacing the original U-Net attention. This mechanism lets the background location actively query and absorb Subject features, yielding a truly semantically fused result that far exceeds vanilla control-map overlays.
Attention-driven Control Fusion
We introduce a self-supervised gating scheme that utilizes the UNet’s own self-attention at each denoising step to produce a raw attention map, thereby enabling the model to autonomously identify regions that warrant additional emphasis. This map is then turned into a crisp neural sketch $\mathcal{S}_{\mathrm{neur}}$ via simple Gaussian smoothing, outlining the key structural skeleton. We fuse this sketch with the original control image in a parameter-free, element-wise manner (e.g., amplifying control features by taking the pixel-wise max), yielding a true attention-driven control signal without any ground-truth supervision.
$I_{\mathrm{fused}}=[\mathcal{S}_{\mathrm{neur}}\vee I_{\mathrm{control}}].$ (8)
When combined with our Cross-branch Attention, this attention-driven fusion reinforces Subject cues within Surrounding regions and refines context coherence, producing sharper strokes and richer textures. And with multiple subjects and concepts, we can extend our framework to multi-concept generation (seen in supplementary materials).
@@ 4 Experiments
Figure 6: Visualization of VitaGlyph’s applications. The experiments show that our method can be applied to other languages and fonts. Additionally, VitaGlyph can be extended to generate with multiple customized concepts and kinetic typography.
Methods Language CLIP Score $\uparrow$ GPT-4o(Aes.) $\uparrow$ GPT-4o(OCR) $\uparrow$ FID Score $\downarrow$
WordArt Designer (He et al. 2023) Chinese 23.93 85.12 88.33% 249.12
VitaGlyph 26.58 94.35 92.23% 236.15
Word-as-image (Iluz et al. 2023) English 29.01 70.66 80.52% 255.28
VitaGlyph 29.16 92.18 81.96% 247.17
@@ 4.1 Implementation Details
Benchmark
We collect two hundred Chinese words (including single and multiple characters) as an evaluation benchmark and ask ChatGPT to annotate subject and surrounding descriptions. Considering that Word-as-image (Iluz et al. 2023) only generates English words, we also translate the collected benchmark into English.
More details are provided in the supplementary materials.
Evaluation Metrics
We use CLIP (Radford et al. 2021) image-text similarity to assess similarities between rendered images and prompt descriptions (i.e., $P_{\mathrm{sub}}$+ $P_{\mathrm{surr}}$) and FID (Heusel et al. 2017) scores to assess readability. Moreover, to better align with human evaluators, we also employ GPT-4o (OpenAI 2024b) to rate the aesthetic scores (GPT-4o Aes.) and determine whether rendered images and can be correctly recognized as input characters (GPT-4o OCR).
@@ 4.2 Quantitative Comparisons
As shown in Tab. 1, we conduct quantitative experiments on the benchmark mentioned above. For comparisons, we leverage WordArt Designer (He et al. 2023) for Chinese words and Word-as-image (Iluz et al. 2023) for English version. Since MetaDesigner has not open-sourced the code and we can only generate results one by one through the website demo. We have only conducted a qualitative comparison with it.
Our VitaGlyph outperforms WordArt Designer in all metrics, especially in OCR with a 3.9% improvement. We believe this is because, with the controllable and proper transformation of glyph geometry, our method maintains promising legibility. Meanwhile, for English words, VitaGlyph also demonstrates strong generalization capabilities, especially in aesthetic field (Aes.) with an improvement of 21.52. This further proves that our method can better express essential concepts with elaborately designed subject transformations.
Since all methods’ output images are in color while the ground-truth images for FID evaluation are black-and-white, all methods result in high FID scores. However, we still outperform other baselines.
@@ 4.3 Qualitative Comparisons
We compare VitaGlyph with WordArt Designer and MetaDesigner, both of which are strong baselines for Chinese artistic typography. We also design English words for a qualitative comparison with Word-as-image. To ensure a fair comparison, we concatenate $P_{\mathrm{sub}}$ and $P_{\mathrm{surr}}$ from our method and use it as input prompt for other works.
@@
In Fig. 5, we see that VitaGlyph renders input characters with higher quality and legibility than WordArt Designer and achieves results that are comparable to, or even better than, those of MetaDesigner. For example, in column $1$, VitaGlyph visualizes the concept of “tomato” more concisely, incorporating related elements such as “fries”. However, the deformation regions in WordArt’s images are not well defined, leading to unnatural distortions and structural damage. And for those from MetaDesigner, the failure in deformation resulted in the fusion of character and background, ultimately transforming the glyph background into a “tomato”.
For the English version, the key concepts for the “Hello World” are designed manually as specialties such as “the Earth, a doughnut, a robot waving its hand and colorful candy texture”. But we can observe the both WordArt and Metadesigner miss the element of “the Earth”. As for Word-as-Image, it excels in structural transformation but it can only generate black-white typography. In contrast, VitaGlyph properly generates imaginable artistic typography with the controllable transformation. More VitaGlyph examples can be found in supplementary materials.
4.4 Ablation Studies
[…]
Aes. $\uparrow$ OCR $\uparrow$
w/o $\texttt{GDINO}_{\mathrm{rank}}$ 23.90 84.47 89.15 255.34
w/o SemTypo 23.53 85.79 92.98 241.58
w/o Attn 25.81 92.74 92.23 239.27
Ours 26.58 94.35 92.23 236.15
Table 2: Ablation of modules in our method. We provide our ablation studies’ quantitative results with different modules removed.
Effect of Region Interpretation
To study the effect of our Region Interpretation from Sec.3.2, we conduct ablation experiments without $\texttt{GDINO}_{\mathrm{rank}}$. Specifically, we regard $I$ as $I_{\mathrm{sub}}$, while keeping the settings for SemTypo unchanged. We solely utilize $\mathrm{controlnet}_{\mathrm{sub}}$ with $I^{*}_{\mathrm{sub}}$ as the additional condition. The input prompt is a concatenation of $P_{\mathrm{sub}}$ and $P_{\mathrm{surr}}$. The results are shown in first row of Fig. 7. Furthermore, we also randomly select the subject region rather than utilize $\texttt{GDINO}_{\mathrm{rank}}$, with all other settings unchanged. The results are presented in second row of Fig. 7.
Fig. 7 shows that without Region Interpretation, VitaGlyph faces challenges: (i)Unreasonable region selection makes it difficult to ensure both prompt alignment and readability simultaneously. For example, the structure of “lavender” is deformed (1st row, sub-fig. a) and the features of “cat” are not accurately generated (2nd row, sub-fig. b). (ii)The overall deformation of the characters is abnormal and some key concepts, such as “fries” in “tomato” (1st row, sub-fig. c), are missing.
For quantitative results, compared to rows 2 and 5 of Tab. 2, the CLIP score improves by 1.91% and the Aes. score increases by 9.88 wtih 3.08% in OCR. This is likely because Grounding-DINO identifies the part most resembling the subject prompt, making glyph geometry transformation easier and more natural.
Effect of Semantic Typography
Although Grounding-DINO identifies the part similar to the subject prompt $P_{\mathrm{sub}}$, Semantic Typography is essential for structural transformation. Ablation experiments without this module show that images fail to capture the character’s essence or cause the Subject to appear fragmented(see Supplementary Fig. D). This is due to the gap between $I_{\mathrm{sub}}$ and real-world objects, which the Depth-to-image module helps bridge.
Based on rows 3 and 5 of Tab. 2, we can conclude that VitaGlyph with Semantic Typography would enrich the semantic expression but sacrifice little readability.
Effect of Attentional Compositional Generation
Figure 8: Visualization of ablation study on ACG module.
To verify the effect of our ACG module, we conduct ablation experiments with noise-based compositional generation. Specifically, we remove the cross-branch attention between Subject and Surrounding and simplify the attention-driven control as image-driven control, with other settings unchanged. The 4th and 5th row of Tab. 2 shows the increase of 1.61 in GPT-4o (Aes.). From the Fig. 8, we can also observe more details when ACG module is added.
We validate our ACG module via an ablation that replaces attention-driven fusion with standard noise-based compositional generation (i.e., we disable cross-branch attention and revert from attention-driven back to image-driven control while keeping all other settings fixed.) As reported in Tab. 2 (rows 4–5), restoring ACG raises the GPT-4o aesthetic score by 1.61. Qualitative examples in Fig. 8 also reveal substantially richer detail when ACG is applied.
5 Application
5.1 Multiple Concepts, Fonts and Languages
In the first row of Fig. 6, we demonstrate VitaGlyph’s effectiveness across various fonts and languages, such as “Bold” in English, “Song Traditional” in Chinese and “Noto Sans” in Korean. VitaGlyph not only generates artistic glyph images, but also preserves the original font styles.
Moreover, VitaGlyph is the first work that can extend a single subject to multiple customized subject concepts. As illustrated in Fig. 6, it designs subjects like “the Earth”, “compass” and “colorful doughnut” within the artistic typography of “Google”. To address the blurred character details in human generation, we integrate VLM (OpenAI 2024b) to drive fine-grained detail enhancement. This plays a crucial role in areas such as commercial advertising, brand identity, and corporate image design.
5.2 Kinetic Artistic Typography
Our framework adapts to kinetic artistic typography by leveraging I2V models (Wan et al. 2025) to animate the Subject alone, enabling controllable, semantics-driven glyph motion. Unlike prior methods that emphasize depth or background-to-glyph transitions (He et al. 2023), we focus on the core motif’s behavior and movement for more expressive animations (Fig. 6).
6 Conclusion
This paper presents VitaGlyph, a novel approach to artistic typography that balances creativity and readability through a controllable glyph transformation. By distinguishing the subject and the surrounding elements, VitaGlyph enhances visual representation while preserving the core meaning of the character. Our method makes significant contributions to artistic text generation. Future work will focus on further advancements in generative AI to enhance the adaptability of typography across diverse contexts and styles. VitaGlyph will significantly enrich artistic expression, enabling more personalized, dynamic designs.
Supplementary Materials
Figure A: More examples of VitaGlyph’s application. Our method can be applied to various languages and fonts. With multiple user-specific concepts injected and the utilization of Attentional Compositional Generation, we manage to design slogan like “Hello World”..
I Prompt Examples
I.1 LLM Templates:
Suppose you are a creative and active explainer, dedicated to helping people understand abstract concepts and concrete them. Your task is to identify and elucidate the representative tangible objects within these abstract concepts, helping the audience better connect abstract thinking with real-world entities. All questions should follow a standardized format, such as: “Please list the representative tangible objects in/of <CONCEPT >, along with the appropriate artistic font style.”Your response must adhere to a strict JSON format, specifically: {“subject prompt”: “representative tangible object that can express the intrinsic semantic of input <CONCEPT >”, “surrounding prompt”: “appropriate artistic font style or texture that can enrich <CONCEPT >.” }This format ensures clarity and consistency in responses, making the information easy to parse and understand. For example, suppose the question is: “Please list the representative category or object name in/of ‘cat’, including in real-life, artist works, and film works, and select an appropriate artistic font style.”Your response should be: {“subject prompt”: “cake and frosting, sprinkles, layers, with features like sweet, colorful, decadent, multi-tiered.”,“surrounding prompt”: “texture design is frosting, sprinkles, layers, with creamy, glossy, textured, and delightful details.”}”
I.2 Subject and Surrounding prompt examples
$Q_{s}$: “Please list the representative tangible objects in/of plum, along with the appropriate artistic font style.”
$A_{s}$: “{“subject prompt”: “round plum, juicy fruit and vibrant blossom with features such as juicy, round, vibrant, fragrant and delicate.”, “surrounding prompt”: “colorful blossom, green leaves, petals, water droplets, velvety skin texture and intricate floral motifs.”}”
$Q_{s}$: “Please list the representative tangible objects in/of rose, along with the appropriate artistic font style.”
$A_{s}$: “{“subject prompt”: “a blooming red rose flower”, “surrounding prompt”: “brown vignette, leaf, slender branch, thorns in a romantic atmosphere”}”
II Noise-based Compositional Generation
The overall Attentional Compositional Generation process is illustrated in Fig. 4. Since the sampling processes for the subject and surrounding are conducted independently, and the U-Net outputs are also independent, directly merging them can result in a visually incoherent image, as seen in Fig.D.
In our method, we resolve this fragmentation issue by allowing the subject and surrounding diffusion processes to share the same unconditional noise $\epsilon^{\mathrm{uc}}_{t}$ during ACG module. $\epsilon^{\mathrm{overall}}_{t}$ can be further harmonized as follows:
where $s$ represents the guidance scale and is set to $7.5$ by default.
III Multi-concept Generation
For $N$ subject concepts, we iteratively detect the subject region for the $i$-th concept with GroundingDINO. This can be formalized as follows:
$I_{\mathrm{sub}}^{i},M^{i}=\texttt{GDINO}_{\mathrm{rank}}(I\setminus\bigcup_{j=1}^{i-1}I_{\mathrm{sub}}^{j},P_{\mathrm{sub}}^{i}),\vspace{-10pt}$ (10)
Then we obtain multiple subject structures and their corresponding masks, allowing us to achieve a multi-concept scenario. The overall predicted noise is computed as:
$\epsilon^{\mathrm{overall}}_{t}=\sum_{i=1}^{N}\gamma^{i}\cdot M^{i}\cdot\epsilon^{\mathrm{sub,i}}_{t}+(1-\sum_{i=1}^{N}M^{i})\cdot\epsilon^{\mathrm{surr}}_{t}.\vspace{-8pt}$ (11)
IV Benchmark Details
Figure B: Benchmark words of our experiments.
We collect 200 Chinese words (including single and multiple characters) as an evaluation benchmark, as shown in Fig .B. For every word, we generate 100 samples and we totally create 20000 images as a dataset. Considering that Word-as-image (Iluz et al. 2023) only generates English words, we also translate the collected benchmark into an English version with 50 samples per word and a total of 10000 images as a dataset.
For words’ descriptions, we ask ChatGPT (OpenAI 2024a) to annotate subject and surrounding prompts. Since other methods, Word-as-image (Iluz et al. 2023), WordArt Designer (He et al. 2023) and MetaDesigner (He et al. 2024) can only leverage one input prompt, we concatenate the subject and surrounding prompts.
V Deployment Details
In phase Knowledge Acquisition, we use GPT3.5 as our Large Language Model, and the output of the LLM is JSON format. In phase Regional Decomposition, we use Grounding-DINO with its hyper-parameters all default to the  (Liu et al. 2023a). After that, we select the bounding boxes with its confidence over 0.5 and its area proportion of the whole image between 0.4 and 0.6 and then choose the highest confidence as the subject part. In phase Typography Stylization, we use Stable Diffusion-XL (Podell et al. 2023) as our base model. All generated images are $512\times 512$ for both the Depth-to-image and ACG module.
VI Effect of LLM’s random outputs
Methods Language CLIP Score $\uparrow$ GPT-4o(Aes.) $\uparrow$ GPT-4o(OCR) $\uparrow$
WordArt Designer (He et al. 2023) Chinese 23.93$\pm$1.55 85.12$\pm$5.96 88.33$\pm$1.95
VitaGlyph 26.58$\pm$1.62 94.34$\pm$4.75 92.23$\pm$1.93
Word-as-image (Iluz et al. 2023) English 29.01$\pm$1.23 70.66$\pm$5.78 80.52$\pm$1.24
VitaGlyph 29.16$\pm$1.28 92.18$\pm$5.26 81.96$\pm$1.79
Table A: Mean & Std of quantitative experiments over five turns with different input prompts.
We perform five regeneration rounds for both $P_{\mathrm{sub}}$ and $P_{\mathrm{surr}}$. The results in Tab. A show that although ChatGPT’s random output actually impacts the final results, its strong comprehension keeps the deviation within an acceptable range.
VII More Ablation Studies
Figure C: Visualization of ablation studies on SemTypo.
VII.1 Effect of shared-$\boldsymbol{\epsilon_{T}}$
Figure D: Visualization of ablation studies on shared-$\boldsymbol{\epsilon_{T}}$.
As VitaGlyph employs two ControlNets to individually render subject and surrounding regions, we investigate the impact of shared initial noise (i.e., shared-$\epsilon_{T}$) on the final results. Fig. D indicates that not using shared-$\epsilon_{T}$ leads to inconsistent appearance and unrealistic lighting. For example, in the first column of “rose” and “monkey”, the colors of the subject and surrounding regions are visibly incoherent.
VII.2 The Strength of Added Noise in SemTypo
Figure E: Ablation Study on D2I’s strength $\mathrm{DS}$ and ControlNet’s scale $\mathrm{CS}$. (a) shows that as $\mathrm{DS}$ increases, the deformation of $I^{*}_{\mathrm{sub}}$ grows during SemTypo. (b) demonstrates the generated image morphology aligns more closely with $I^{*}_{\mathrm{sub}}$ as $\mathrm{CS}$ increases.
Semantic Typography uses the SDEdit algorithm (Meng et al. 2021) based on the Depth-to-image diffusion model to adjust the geometry of the subject region. As shown in Fig. E (a), as the strength of added noise ($\mathrm{DS}$) increases, more structural details are added to the subject region while it remains in black-and-white color. When $\mathrm{DS}$ reaches $0.85$, the subject region contains enough details of the given concept, e.g., the ears and tails of cats in the first row.
VII.3 The Conditioning Scale in ControlNet
We fix $I^{*}_{\mathrm{sub}}$ and only adjust the $\mathrm{ControlNet}_{\mathrm{sub}}$’s guidance scale $\mathrm{CS}$. From Fig. E (b), we observe that as the $\mathrm{CS}$ increases, the generated part of Subject becomes more clear and more similar to $I^{*}_{\mathrm{sub}}$, shown in row 1. However, being too similar to the subject can lead to loss of detail, as seen in row 2 where the flower’s stamen disappears from $\mathrm{CS}=1.0$.
VIII User Study
Methods Text OCR Aesthetic
@@ VIII User Study
Table B: User Preference Study. Thirty users participated in the survey, selecting images that best represent the concept (Text Align.), match the character (OCR), and show the highest artistic mastery (Aesthetic).
@@
We conduct a user study to evaluate our method against WordArt Designer and Word-as-Image on 80 samples, each consisting of an input character and three images from VitaGlyph, WordArt and Word-as-image (random order). For each sample, we ask 30 workers to select the image that best represents the concept, matches the character, and is the most aesthetically pleasing, yielding 2400 total responses. The statistical results in Tab. B show that users prefer the rendered images from VitaGlyph across all metrics.
IX Additional Results